# From Fog to Failure: How Dehazing Can Harm Clear Image Object Detection

### v1

@@ A.1 Foggy CityScapes
The Foggy Cityscapes (Sakaridis et al., 2018) dataset is created to address the problem of semantic foggy scene understanding (SFSU). While there has been extensive research on image dehazing and semantic scene understanding with clear-weather images, SFSU has received little attention. Due to the difficulty in collecting and annotating foggy images, synthetic fog is added to real images depicting clear-weather outdoor scenes. This synthetic fog generation leverages incomplete depth information to create realistic foggy conditions on images from the Cityscapes dataset, resulting in Foggy Cityscapes with 20,550 images. The training set consists of 2975 images, validated on 500 images, and the test set had 1525 images. The key features of the dataset include:
  • Synthetic Fog Generation: Real clear-weather images are used, and synthetic fog is added using a complete pipeline that employs the transmission map.
  • Data Utilization: The dataset can be utilized for supervised learning and semi-supervised learning. We generated a foggy dataset using the synthetic transmission map and then performed supervised learning on the synthetic foggy data.
@@ A.2 RESIDE-$\beta$
The RESIDE-$\beta$ Outdoor Training Set (OTS) is a comprehensive dataset designed to facilitate research in outdoor image dehazing. It addresses the challenges posed by haze in outdoor scenes, which significantly degrades image quality and affects subsequent tasks such as object detection and semantic segmentation. The dataset includes approximately 72,135 outdoor images with varying degrees of haze, enabling robust training of dehazing algorithms. For testing, we are using the RESIDE-$\beta$ (REalistic Single Image DEhazing) dataset (Li et al., 2019). The subset of RESIDE-$\beta$, Real-Time Testing Set (RTTS) consists of 4,322 real-world hazy images with annotations for object detection. The training set consists of 3000 images, validated on 500 images, and the test set had 1500 images.
@@ Appendix B Human Visual Cues
Selective Attention and Foveation: The human eye is not equally sensitive to all parts of the visual field. Central vision, or foveal vision, is highly detailed and used for tasks like reading and identifying objects. Peripheral vision is less detailed and more sensitive to motion. The system scans the entire image (peripheral vision) similar to our preliminary detection phase. This will identify areas for a more detailed analysis (foveal vision), mimicking the human approach of not processing every detail with equal clarity but focusing on areas of interest.
Adaptation to environmental conditions: Just as the human visual system adapts to different lighting conditions and levels of visibility (such as adjusting to a dark room after being in bright sunlight), the adaptive dehazing method adjusts the intensity and focus of its processing based on the detection feedback and environmental context, analogous to the way human vision adjusts to ensure optimal perception under varying conditions.
Eye Tracking and Gaze-directed processing: Eye-tracking monitors where a person is looking (the gaze) and what draws attention. In visual processing, this is analogous to directing computational resources toward areas of interest, much like the proposed method focuses on dehazing and detailed detection of regions where objects are likely to be present. By analogy, the system pays more attention to certain parts of the image, just as a person would fixate on specific areas within their field of view when searching for something.
Integration of Bottom-up and Top-down processes: The human visual system uses both bottom-up signals (from sensory input) and top-down processes (based on knowledge, expectations, and current goals) to interpret scenes. The proposed model initially uses a bottom-up approach (object detection algorithms flagging potential areas of interest) followed by a top-down approach (focusing on dehazing efforts based on three flagged areas and previous learning), mirroring the complex interplay between sensory data and cognitive processes in human vision.
@@ Appendix C Methodology
Preliminary Detection: Implement a lightweight, fast object detection algorithm such as YOLOv5s or YOLOv8n to quickly scan the image for potential regions of interest or active regions and flag those patches with a high likelihood of containing objects. The smaller versions of standard YOLO models are less accurate than their full-sized counterparts but significantly faster, making them ideal for preliminary detection.
Region-based dehazing: Apply dehazing algorithms specifically to those active regions identified in the preliminary detection phase. Considering the depth or level of haze, the method should adapt based on the characteristics of the detected regions
The proposed architecture of AOD-NetX in Figure 2 utilizes the transmission map created by the standard AOD-Net (Li et al., 2017a) and applies it within a spatial attention map module to produce an attention-focused transmission map. This spatial attention map is derived from the bounding boxes or Regions of Interest identified by the lightweight model (YOLOv5s/YOLOv8n) in our proposed method. A sigmoid layer follows, mapping the output probabilities to a range between 0 and 1. We opt not to use softmax in this context due to the independent significance of each bounding box.
@@ D.1 AOD-Net
AOD-Net, or All-in-One Dehazing Network, is a convolutional neural network (CNN) designed to remove haze from images by directly reconstructing the clean image through an end-to-end approach. Unlike traditional methods that separately estimate transmission maps and atmospheric light, AOD-Net is based on a re-formulated atmospheric scattering model, enabling it to generate dehazed images without intermediate estimations. This lightweight architecture not only achieves superior performance in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) but also enhances visual quality. Moreover, its design allows for easy integration into other deep learning models, such as Faster R-CNN, improving performance on tasks like object detection in hazy conditions.
@@ D.2 Dehaze-UNet
The U-Net dehaze technique uses an encoder-decoder structure with skip connections for effective image dehazing. In the encoder, the network reduces spatial dimensions through double convolution layers, extracting essential features. Each convolution is followed by batch normalization and ReLU activation to enhance feature representation. In the decoder, these features are upsampled, reintegrating spatial details from the encoder via skip connections, crucial for restoring fine details lost in downsampling. Double convolution layers in the decoder refine the combined features, ensuring clarity. The model concludes with Max Pooling and Bilinear Interpolation, optimizing feature representation and smoothing transitions in the upsampled images, making U-Net highly effective in dehazing tasks.
@@ D.3 DehazeNet
The DehazeNet technique has four sequential steps presented in Figure 5 [35]. The feature extraction combines 16 filters with 4 MaxOut layers that perform max pooling to reduce data dimensionality [35]. Multi-scale mapping ensures that features of different complexity (both detailed and high-level features) are considered. Local extremum acts as max pooling that is specifically tuned to achieve spatial invariance, preserving the image resolution. In the final stage of Non-linear regression, the Bilateral ReLU is used as an activation function; it bounds the output within a specific range to avoid oversaturated regions
@@ D.4 AOD-NetX
The AOD-NetX architecture as shown in Figure 2 utilizes the transmission map created by the standard AOD-Net and applies it within a spatial attention map module to produce an attention-focused transmission map. This spatial attention map is derived from the bounding boxes or Regions of Interest identified by the lightweight model (YOLOv5s) in our proposed method. A sigmoid layer follows, mapping the output probabilities to a range between 0 and 1. We opt not to use softmax in this context due to the independent significance of each bounding box.
@@ Appendix E Object Detection Models
The YOLO models used in the detection pipeline include a variety of versions optimized for different purposes. YOLOv5s is a lightweight variant designed for real-time detection with low computational requirements, while YOLOv8n (Nano) is tailored for high-speed applications on resource-constrained devices like mobile phones. On the other hand, YOLOv5x, with its CSP backbone and advanced data augmentation, provides enhanced performance for more complex scenes, and YOLOv8x (Extra Large) offers maximum accuracy for large-scale datasets. The detection process begins by using YOLOv5s or YOLOv8n on foggy images to generate initial annotations. These annotations, along with the original image, are then dehazed using AOD-NetX, and the dehazed image is subsequently passed through YOLOv5x or YOLOv8x for precise and refined detection results.
@@ Appendix F Additional Results
The dehazing modules are trained separately on the provided datasets, while the object detection models (various YOLO versions) remain pre-trained on the MS-COCO dataset. This approach allows users to integrate the dehazing module with their own detection pipeline without requiring a complete re-training of the entire system. However, for improved results, the entire architecture could be fine-tuned on the target datasets, which would serve as a valuable direction for future ablation studies.
The results in Table 3 show that AOD-NetX generally outperforms the standard AOD-Net in terms of SSIM and PSNR across most datasets. For Foggy Cityscapes and RESIDE-$\beta$ OTS, AOD-NetX achieves higher SSIM and PSNR, indicating improved structural similarity and signal quality. However, for RESIDE-$\beta$ RTTS, while AOD-NetX has a slightly better PSNR, AOD-Net achieves a significantly higher SSIM score, suggesting that AOD-Net may retain more structural details in this particular dataset. Overall, AOD-NetX is more effective in most scenarios, especially for complex foggy conditions.
@@ F.1 In-distribution performance of Perceptual Piercing
The evaluation results of Perceptual Piercing variations trained and tested on the Foggy Cityscapes dataset has been provided in Table 4. The integration of dehazing modules, such as AOD-Net and AOD-NetX (detailed in Appendix C), consistently improves object detection in both clear and foggy conditions. The AOD-Net + YOLOv5x variant achieved the highest mAP under clear conditions (0.6813), while YOLOv5s + AOD-NetX + YOLOv5x and YOLOv8n + AOD-NetX + YOLOv8x performed best in foggy conditions, with mAP scores of 0.6152 and 0.6114, respectively. In contrast, baseline YOLO models (YOLOv5x and YOLOv8x) exhibited lower detection accuracy, underscoring the effectiveness of enhanced dehazing techniques in low-visibility environments.
@@ F.2 Out-of-distribution performance of Perceptual Piercing
The evaluation of various Perceptual Piercing variations in Table 5 trained on Foggy Cityscapes and tested on RESIDE-$\beta$ OTS and RTTS datasets shows that the YOLOv8x architecture achieved the highest mAP scores under foggy conditions, with 0.7125 on OTS and 0.6978 on RTTS. Among the YOLOv5 variants, the baseline YOLOv5x model performed the best, with 0.6944 on OTS and 0.6655 on RTTS. The addition of AOD-Net generally improved performance for YOLOv8 but had a diminishing effect on YOLOv5. Models incorporating AOD-NetX showed lower mAP values across both test datasets, indicating that its integration may need further optimization. Overall, the results suggest that YOLOv8x is more robust for foggy conditions compared to other variations.
@@ G.1 Structural Similarity Index Measure (SSIM)
The performance of dehazing methods is evaluated by the following equation of SSIM score between two images:
@@
  • $\mu_{x}$ and $\mu_{y}$ are the average of $x$ and $y$ respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the variance of $x$ and $y$ respectively.
  • $\sigma_{xy}$ is the covariance of $x$ and $y$.
  • $c_{1}=(k_{1}L)^{2}$ and $c_{2}=(k_{2}L)^{2}$ are two variables to stabilize the division with weak denominator; $L$ is the dynamic range of the pixel-values (typically this is $12^{bits\_per\_pixel}-$ 1), $k_{1}=0.01$ and $k_{2}=0.03$ by default.
@@ G.2 Peak Signal-to-Noise ratio (PSNR)
Peak Signal-to-Noise Ratio (PSNR) is a widely used metric for evaluating the quality of reconstructed images or videos compared to the original, reference data. It is expressed in decibels (dB) and is calculated based on the mean squared error (MSE) between the original and the reconstructed images. The formula for PSNR is given by:
@@
where MAX is the maximum possible pixel value of the image (for example, 255 for 8-bit images), and MSE is the mean squared error, defined as:
$\text{MSE}=\frac{1}{mn}\sum_{i=0}^{m-1}\sum_{j=0}^{n-1}\left(I(i,j)-K(i,j)\right)^{2},$ (3)
where $I(i,j)$ represents the pixel value at position $(i,j)$ in the original image, and $K(i,j)$ represents the pixel value at the same position in the reconstructed image. Higher PSNR values generally indicate better reconstruction quality, as they imply a lower MSE and thus less distortion. PSNR is particularly useful for comparing the performance of different image processing algorithms in tasks such as image compression, denoising, and super-resolution.
G.2.1 Mean Average Precision (mAP)
For object detection performance, we are using mean Average Precision (mAP):
$AP=\frac{\sum_{k=1}^{n}(P(k)\times\text{rel}(k))}{\text{number of relevant documents}}$ (4)
@@ G.2.1 Mean Average Precision (mAP)
  • $P(k)$ is the precision at cutoff $k$ in the list.
  • $\text{rel}(k)$ is an indicator function equaling 1 if the item at rank $k$ is a relevant document, 0 otherwise.
  • $n$ is the number of retrieved documents.
The mean Average Precision is then calculated as:
@@
where $AP_{q}$ is the Average Precision for the $q^{th}$ query and $Q$ is the total number of queries.

### v2

@@ A.1 Foggy CityScapes
The Foggy Cityscapes dataset (Sakaridis et al., 2018) was developed to tackle the challenge of semantic foggy scene understanding (SFSU). While significant research has been conducted on image dehazing and semantic scene understanding for clear-weather images, SFSU remains relatively underexplored. Due to the challenges associated with collecting and annotating real-world foggy images, synthetic fog is introduced into clear-weather outdoor scenes. This synthetic fog generation process utilizes incomplete depth information to simulate realistic foggy conditions on images from the Cityscapes dataset, resulting in a dataset comprising 20,550 images. The dataset is divided into a training set of 2,975 images, a validation set of 500 images, and a test set of 1,525 images. Key characteristics of the dataset include:
  • Synthetic Fog Generation: Synthetic fog is added to real clear-weather images using a dedicated pipeline that incorporates the transmission map.
  • Data Utilization: The dataset supports both supervised and semi-supervised learning. A synthetic foggy dataset was generated using the synthetic transmission map, followed by supervised learning on the resulting foggy images.
@@ A.2 RESIDE-$\beta$
The RESIDE-$\beta$ Outdoor Training Set (OTS) is a comprehensive dataset curated to support research in outdoor image dehazing. It addresses the degradation caused by haze in outdoor scenes, which negatively impacts image quality and downstream tasks such as object detection and semantic segmentation. The dataset contains approximately 72,135 outdoor images with varying haze intensities, allowing for robust training of dehazing algorithms. For evaluation, we use the RESIDE-$\beta$ (REalistic Single Image DEhazing) dataset (Li et al., 2019). A subset of RESIDE-$\beta$ , the Real-Time Testing Set (RTTS), comprises 4,322 real-world hazy images with object detection annotations. The dataset is split into a training set of 3,000 images, a validation set of 500 images, and a test set of 1,500 images.
@@ Appendix B Human Visual Cues
Selective Attention and Foveation: The human eye does not perceive all areas of the visual field with equal clarity. Foveal vision, which corresponds to central vision, is highly detailed and is essential for tasks such as reading and object recognition. In contrast, peripheral vision is less detailed but more sensitive to motion. The visual system initially scans the entire scene using peripheral vision, akin to the preliminary detection phase in our approach. This broad scanning process helps identify regions requiring closer inspection, enabling a more detailed analysis through foveal vision. Similarly, the proposed method does not process every detail uniformly but prioritizes key areas of interest.
Adaptation to Environmental Conditions: The human visual system dynamically adjusts to varying lighting conditions and levels of visibility, such as adapting from bright sunlight to a dark room. Similarly, the adaptive dehazing method modulates its processing intensity and focus based on detection feedback and environmental context. This mechanism ensures optimal perception, mirroring the way human vision adapts to maintain clarity under diverse conditions.
Eye Tracking and Gaze-Directed Processing: Eye-tracking technology monitors gaze direction and identifies focal points of attention. This concept translates to strategically allocating resources toward regions of interest in computational visual processing. The proposed method follows a similar principle by directing dehazing and detailed object detection efforts to areas where objects are likely to be present. Just as human vision selectively fixates on specific regions when searching for an object, the system prioritizes certain parts of the image to enhance clarity and detection performance.
Integration of Bottom-Up and Top-Down Processes: Human vision combines bottom-up processing, driven by sensory input, with top-down processing, guided by prior knowledge, expectations, and goals. The proposed model adopts a similar dual approach: it first employs a bottom-up strategy, where object detection algorithms identify potential areas of interest. This is followed by a top-down refinement process, where dehazing efforts are concentrated on flagged areas, leveraging previous learning. This interplay between data-driven signals and cognitive insights aligns with the way human perception integrates sensory input with contextual understanding.
@@ Appendix C Methodology
Preliminary Detection: A lightweight and fast object detection algorithm, such as YOLOv5s or YOLOv8n, is employed to rapidly scan the image and identify potential regions of interest or active regions. These models flag image patches with a high probability of containing objects. While the smaller variants of YOLO models offer lower accuracy compared to their full-sized counterparts, they are significantly faster, making them well-suited for this initial detection phase.
Region-Based Dehazing: Dehazing algorithms are selectively applied to the active regions identified during the preliminary detection phase. The approach dynamically adjusts based on the depth or severity of haze within the detected regions, ensuring an adaptive and efficient dehazing process.
The proposed architecture, AOD-NetX, illustrated in Figure 2, builds upon the transmission map generated by the standard AOD-Net (Li et al., 2017a). This transmission map is integrated into a spatial attention map module, producing an attention-enhanced transmission map. The spatial attention map is derived from the bounding boxes or Regions of Interest (ROIs) detected by the lightweight object detection model (YOLOv5s/YOLOv8n) within the proposed framework. A sigmoid layer is applied to map the output probabilities to a range between 0 and 1. Unlike softmax, which normalizes outputs across multiple regions, sigmoid is preferred in this context since each bounding box holds independent significance.
@@ D.1 AOD-Net
AOD-Net (All-in-One Dehazing Network) is a convolutional neural network (CNN) designed for haze removal by directly reconstructing the clean image in an end-to-end manner. Unlike traditional approaches that separately estimate transmission maps and atmospheric light, AOD-Net is based on a re-formulated atmospheric scattering model, allowing it to generate dehazed images without intermediate computations. This lightweight architecture delivers superior performance in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) while also enhancing visual quality. Additionally, its modular design enables seamless integration into other deep learning models, such as Faster R-CNN, thereby improving object detection performance in hazy conditions.
@@ D.2 Dehaze-UNet
The Dehaze-UNet model employs an encoder-decoder architecture with skip connections to effectively restore dehazed images. The encoder reduces spatial dimensions through successive convolutional layers, extracting essential features while maintaining structural integrity. Each convolutional layer is followed by batch normalization and ReLU activation, enhancing feature representation. The decoder then upscales these features, reintegrating spatial details from the encoder via skip connections, which are critical for preserving fine-grained textures lost during downsampling. The final stage consists of double convolution layers in the decoder for feature refinement, followed by Max Pooling and Bilinear Interpolation to optimize feature representation and ensure smooth transitions in the reconstructed image. This makes U-Net particularly effective for dehazing tasks.
@@ D.3 DehazeNet
DehazeNet follows a structured four-stage approach, as illustrated in Figure 5 (Zhang & Patel, 2018). The first stage, feature extraction, employs 16 filters combined with four MaxOut layers, which use max pooling to reduce data dimensionality. The second stage, multi-scale mapping, captures both fine-grained and high-level features to ensure a comprehensive feature representation. The third stage, local extremum, utilizes a specialized max pooling operation to enhance spatial invariance while preserving image resolution. Finally, the non-linear regression stage incorporates Bilateral ReLU as the activation function, which constrains the output within a defined range to prevent oversaturation and maintain image clarity.
@@ D.4 AOD-NetX
The proposed AOD-NetX architecture, depicted in Figure 2, extends AOD-Net by leveraging its transmission map within a spatial attention map module to generate an attention-focused transmission map. This spatial attention map is derived from the bounding boxes or Regions of Interest (ROIs) identified by the lightweight object detection model (YOLOv5s) within our framework. A sigmoid activation layer is applied to map the output probabilities to a range between 0 and 1. Unlike softmax, which normalizes outputs across multiple regions, sigmoid is preferred as each bounding box holds independent significance.
@@ Appendix E Object Detection Models
The detection pipeline incorporates various YOLO models, each optimized for specific applications. YOLOv5s is a lightweight variant designed for real-time detection with minimal computational overhead, while YOLOv8n (Nano) is optimized for high-speed processing on resource-constrained devices, such as mobile phones. In contrast, YOLOv5x, with its CSP backbone and advanced data augmentation techniques, delivers enhanced performance for more complex scenes, whereas YOLOv8x (Extra Large) achieves maximum accuracy when handling large-scale datasets.
The detection workflow begins by applying YOLOv5s or YOLOv8n to foggy images to generate initial object annotations. These annotations, along with the original image, undergo dehazing using AOD-NetX. The resulting dehazed image is then processed with YOLOv5x or YOLOv8x, ensuring precise and refined detection outcomes.
@@ Appendix F Additional Results
The dehazing modules are trained independently on the provided datasets, while the object detection models (various YOLO versions) remain pre-trained on the MS-COCO dataset. This modular approach allows seamless integration of the dehazing module into existing detection pipelines without requiring full retraining. However, fine-tuning the entire architecture on target datasets could yield further performance improvements, making it a promising direction for future ablation studies.
Table 3 presents the comparative results, demonstrating that AOD-NetX generally outperforms the standard AOD-Net in terms of SSIM and PSNR across most datasets. For Foggy Cityscapes and RESIDE-$\beta$ OTS, AOD-NetX achieves higher SSIM and PSNR values, indicating superior structural similarity and signal fidelity. However, in the case of RESIDE-$\beta$ RTTS, while AOD-NetX attains a slightly higher PSNR, AOD-Net exhibits a significantly higher SSIM score, suggesting better structural detail retention in this specific dataset. Overall, AOD-NetX proves more effective in most scenarios, particularly under complex foggy conditions.
@@ F.1 In-Distribution Performance of Perceptual Piercing
The evaluation results of Perceptual Piercing variations, trained and tested on the Foggy Cityscapes dataset, are presented in Table 4. The integration of dehazing modules, such as AOD-Net and AOD-NetX (detailed in Appendix C), consistently enhances object detection in both clear and foggy conditions.
Among the tested variants, the AOD-Net + YOLOv5x configuration achieved the highest mAP under clear conditions (0.6813). In foggy conditions, YOLOv5s + AOD-NetX + YOLOv5x and YOLOv8n + AOD-NetX + YOLOv8x demonstrated the best performance, with mAP scores of 0.6152 and 0.6114, respectively. In contrast, baseline YOLO models (YOLOv5x and YOLOv8x) exhibited lower detection accuracy, highlighting the effectiveness of advanced dehazing techniques in low-visibility environments.
@@ F.2 Out-of-Distribution Performance of Perceptual Piercing
The evaluation results in Table 5, where Perceptual Piercing variations were trained on Foggy Cityscapes and tested on the RESIDE-$\beta$ OTS and RTTS datasets, highlight key performance trends. The YOLOv8x architecture achieved the highest mAP scores under foggy conditions, with 0.7125 on OTS and 0.6978 on RTTS. Among the YOLOv5 variants, the baseline YOLOv5x model performed best, achieving 0.6944 on OTS and 0.6655 on RTTS.
The addition of AOD-Net generally enhanced performance for YOLOv8 but had a diminishing effect on YOLOv5. Meanwhile, models incorporating AOD-NetX exhibited lower mAP values across both test datasets, suggesting that its integration may require further optimization. Overall, the results indicate that YOLOv8x is more robust in handling foggy conditions compared to other model variations.
@@ G.1 Structural Similarity Index Measure (SSIM)
The performance of dehazing methods is evaluated using the Structural Similarity Index Measure (SSIM), which quantifies the similarity between two images based on luminance, contrast, and structural components. It is defined as:
@@
  • $\mu_{x}$ and $\mu_{y}$ are the mean intensities of images $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ denote the variances of $x$ and $y$.
  • $\sigma_{xy}$ represents the covariance between $x$ and $y$.
  • $c_{1}=(k_{1}L)^{2}$ and $c_{2}=(k_{2}L)^{2}$ are stabilizing constants to prevent division by zero, where $L$ is the dynamic range of pixel values (e.g., 255 for 8-bit images), and default values are $k_{1}=0.01$ and $k_{2}=0.03$.
@@ G.2 Peak Signal-to-Noise Ratio (PSNR)
Peak Signal-to-Noise Ratio (PSNR) is a widely used metric to assess image reconstruction quality by comparing the original and processed images. It is expressed in decibels (dB) and is calculated as:
@@
where MAX is the maximum possible pixel value (e.g., 255 for 8-bit images), and MSE is the Mean Squared Error:
$\text{MSE}=\frac{1}{mn}\sum_{i=0}^{m-1}\sum_{j=0}^{n-1}\left(I(i,j)-K(i,j)\right)^{2}.$ (3)
Here, $I(i,j)$ and $K(i,j)$ represent pixel values at position $(i,j)$ in the original and reconstructed images, respectively. A higher PSNR value indicates better image quality, as it corresponds to lower distortion. PSNR is extensively used in evaluating dehazing, denoising, and image compression methods.
G.3 Mean Average Precision (mAP)
For object detection performance, we use mean Average Precision (mAP), which evaluates the precision-recall tradeoff. The Average Precision (AP) is computed as:
$AP=\frac{\sum_{k=1}^{n}(P(k)\times\text{rel}(k))}{\text{number of relevant objects}}$ (4)
@@ G.3 Mean Average Precision (mAP)
  • $P(k)$ is the precision at rank $k$.
  • $\text{rel}(k)$ is an indicator function, which is 1 if the object at rank $k$ is relevant, and 0 otherwise.
  • $n$ is the total number of retrieved objects.
The mean Average Precision is computed as:
@@
where $AP_{q}$ is the Average Precision for the $q^{th}$ query, and $Q$ is the total number of queries. Higher mAP values indicate better object detection performance across different classes.